Consider all possible samples of size 2 that can be drawn with replacement

from the population consists of numbers 1,2, and 3


a. mean of the sampling distribution

b. variance of the sampling distribution

c. standard deviation of the sampling distribution

 


1
Expert's answer
2021-05-11T14:14:47-0400

The sampling distribution is shown in the figure-




Total sample N=9


(a) Mean "\bar{x} = \dfrac{ \sum x}{N}=\dfrac{18}{9}=2"


(b) Variance "\sigma^2=\dfrac{\sum (x-\bar{x})^2}{N}=\dfrac{3}{9}=0.33"


(c) Standard deviation "=\sqrt{\sigma^2}=\sqrt{0.33}=0.5773"